# Detection Made Easy: Potentials of Large Language Models for Solidity Vulnerabilities

### v1

@@ Abstract
The large-scale deployment of Solidity smart contracts on the Ethereum mainnet has increasingly attracted financially-motivated attackers in recent years. A few now-infamous attacks in Ethereum’s history includes DAO attack in 2016 (50 million dollars lost), Parity Wallet hack in 2017 (146 million dollars locked), Beautychain’s token BEC in 2018 (900 million dollars market value fell to 0), and NFT gaming blockchain breach in 2022 (\$600 million in Ether stolen). This paper presents a comprehensive investigation of the use of large language models (LLMs) and their capabilities in detecting OWASP Top Ten vulnerabilities in Solidity. We introduce a novel, class-balanced, structured, and labeled dataset named VulSmart, which we use to benchmark and compare the performance of open-source LLMs such as CodeLlama, Llama2, CodeT5 and Falcon, alongside closed-source models like GPT-3.5 Turbo and GPT-4o Mini. Our proposed SmartVD framework is rigorously tested against these models through extensive automated and manual evaluations, utilizing BLEU and ROUGE metrics to assess the effectiveness of vulnerability detection in smart contracts. We also explore three distinct prompting strategies—zero-shot, few-shot, and chain-of-thought—to evaluate the multi-class classification and generative capabilities of the SmartVD framework. Our findings reveal that SmartVD outperforms its open-source counterparts and even exceeds the performance of closed-source base models like GPT-3.5 and GPT-4 Mini. After fine-tuning, the closed-source models, GPT-3.5 Turbo and GPT-4o Mini, achieved remarkable performance with 99% accuracy in detecting vulnerabilities, 94% in identifying their types, and 98% in determining severity. Notably, SmartVD performs best with the ‘chain-of-thought’ prompting technique, whereas the fine-tuned closed-source models excel with the ‘zero-shot’ prompting approach.
@@ I Introduction
This paper presents SmartVD, an efficient LLM-driven vulnerability detection framework for Solidity codes, which is developed by fine-tuning the pre-trained LLM, Codelama. To facilitate this fine-tuning, we introduce a carefully curated domain-specific class-balanced dataset, called VulSmart. Our approach involves two key methods to enable a detailed and accurate assessment by the model: a) binary classification, which determines whether vulnerabilities exist within functions, and b) multi-class classification and generation, which identifies specific types of vulnerabilities along with its severity. This evaluative methodology is consistently applied to both traditional deep learning classification techniques and other popular LLMs under review, enabling us to perform a rigorous assessment on the effectiveness of SmartVD in comparison to other pre-trained LLMs. We also attempt to understand how fine-tuning the LLMs on our labeled datasets affects the model’s performance in this classification task. Finally, we conduct a manual review and alteration of selected code examples to uncover which features LLMs can and cannot recognize when predicting vulnerabilities.
@@ I-A Research Questions
RQ1. How do various pre-trained LLMs and traditional deep learning techniques perform in detecting security vulnerabilities of smart contracts?
RQ2. How do different pre-trained LLMs perform in detecting security vulnerabilities of smart contracts after fine-tuning on VulSmart dataset?
RQ3. How do different prompting techniques impact the efficacy of pre-trained large language models (LLMs) compared to fine-tuned LLMs on the VulSmart dataset?
RQ4: Can adversarial attacks impact the performance of fine-tuned LLMs in detecting vulnerabilities?
@@ II Related Studies
The following work have been relevant to the following three research areas, namely smart contract vulnerability detection using: a) Formal method tools b) Deep learning based approaches, and c) Large Language Models (LLMs)
@@ II-A Formal Method Tools
II-B Deep Learning based methods
Deep learning-based proposals are not dependent on fixed rules for bug detection and are more extendable than formal method tools. There have been a number of proposals to detect vulnerability in smart contracts using various deep learning techniques. In [27] authors employs long short-term memory (LSTM) to sequentially learn smart contract weakness through ethereum opcodes. Liao et al. in [28] introduce soliaudit tool which uses machine learning technique in combination with fuzz testing of smart contracts to detect vulnerability. In [29] authors use average stochastic gradient descent weight-dropped LSTM (AWD-LSTM), a variant of LSTM to detect vulnerability in smart contract. In [30] the control and data flow information of contract is used by graph neural networks and expert knowledge for detecting vulnerability. The authors in [31] introduces Contractward that uses different machine learning techniques that can be used to detect six types of smart contract vulnerabilities. Liu et al. [32] use graph neural network technique to built AMEVuldetector which uses semantic graph feature from source code and local expert patterns techniques to detect vulnerabilities. In [33] the authors propose a deep learning based framework (DeeSCVHunter) in combination with vulnerability candidate slicing to detect reentrancy and time dependence vulnerability in smart contracts. The authors in [34] propose VSCL framework which uses smart contract bytecode represented as feature vector which is fed to deep neural network to detect vulnerability present. Vulpedia [35] uses structural program features from smart contracts along with detection rules composed of vulnerability signatures to detect vulnerability.
II-C Large Language Models
@@ V-D Observations
RQ1. How do various pre-trained LLMs and traditional deep learning techniques perform in detecting security vulnerabilities of smart contracts?
@@
RQ2. How do different pre-trained LLMs perform in detecting security vulnerabilities of smart contracts after fine-tuning on VulSmart dataset?
@@
RQ3. How do different prompting techniques impact the efficacy of pre-trained large language models (LLMs) compared to fine-tuned LLMs on the VulSmart dataset?
@@
RQ4: Can adversarial attacks impact the performance of fine-tuned LLMs in detecting vulnerabilities?
@@ VII Conclusion and future scope
This paper presents a comprehensive approach to leveraging the power of large language models (LLMs) for detecting vulnerabilities in Solidity-based smart contracts. While much of the existing research highlights the potential of closed-source GPT models and Llama2, we introduce the Smart Contract Vulnerability Detection Framework (SmartVD), which demonstrates superior performance in vulnerability detection. Our findings indicate that among open-source LLMs, the fine-tuned CodeLlama (SmartVD) outperforms and along with it perform better than closed-source base models GPT-3.5 and GPT-4o Mini. On the other hand, after fine-tuning, the closed-source GPT-3.5 Turbo and GPT-4o Mini models achieve the highest performance across both open and closed-source LLMs.
@@
Although this work adapts multi-class classification based on the type of vulnerabilities present in the code and their severity levels, our future plan is to strengthen this by pinpointing the exact line or vulnerable function within the code. We also plan to explore the impact on LLM learning by integrating semantic information such as Abstract Syntax Trees (AST), Control Flow Graphs (CFG), and Program Dependency Graphs (PDG) alongside the source code. This integration has significant potential for advancing vulnerability detection tasks.

### v2

@@ Abstract
The large-scale deployment of Solidity smart contracts on the Ethereum mainnet has increasingly attracted financially-motivated attackers in recent years. A few now-infamous attacks in Ethereum’s history includes DAO attack in 2016 (\$50 million lost), Parity Wallet hack in 2017 (\$146 million locked), Beautychain’s token BEC in 2018 (\$900 million market value fell to 0), and NFT gaming blockchain breach in 2022 (\$600 million stolen). This paper presents a comprehensive investigation of the use of large language models (LLMs) and their capabilities in detecting OWASP Top Ten vulnerabilities in Solidity. We introduce a novel, class-balanced, structured, and labeled dataset named VulSmart, which we use to benchmark and compare the performance of open-source LLMs such as CodeLlama, Llama2, CodeT5 and Falcon, alongside closed-source models like GPT-3.5 Turbo and GPT-4o Mini. Our proposed SmartVD framework is rigorously tested against these models through extensive automated and manual evaluations, utilizing BLEU and ROUGE metrics to assess the effectiveness of vulnerability detection in smart contracts. We also explore three distinct prompting strategies—zero-shot, few-shot, and chain-of-thought—to evaluate the multi-class classification and generative capabilities of the SmartVD. Our findings reveal that SmartVD outperforms its open-source counterparts and even exceeds the performance of closed-source base models like GPT-3.5 and GPT-4o Mini. After fine-tuning, the closed-source models, GPT-3.5 Turbo and GPT-4o Mini, achieved remarkable performance with 99% accuracy in detecting vulnerabilities, 94% in identifying their types, and 98% in determining severity. Notably, SmartVD performs best with the ‘chain-of-thought’ prompting technique, whereas the fine-tuned closed-source models excel with the ‘zero-shot’ prompting.
Index Terms:
Large Language Models, Fine-tuning, Solidity, Smart Contracts, Vulnerability Detection.
@@ I Introduction
This paper presents SmartVD, an efficient LLM-driven vulnerability detection framework for Solidity codes, which is developed by fine-tuning the pre-trained LLM, Codelama. To facilitate this fine-tuning, we introduce a carefully curated domain-specific class-balanced dataset, called VulSmart. Our approach involves two key methods to enable a detailed and accurate assessment by the model: a) binary classification, which determines whether vulnerabilities exist within functions, and b) multi-class classification and generation, which identifies specific types of vulnerabilities along with its severity. This evaluative methodology is consistently applied to both traditional deep learning classification techniques and other popular LLMs under review, enabling us to perform a rigorous assessment on the effectiveness of SmartVD in comparison to other pre-trained LLMs. We also attempt to understand how fine-tuning the LLMs on our labeled datasets affects the model’s performance in this classification task. Additionally, we conduct a manual review and alteration of selected code examples to uncover which features LLMs can and cannot recognize when predicting vulnerabilities. Finally, we incorporate semantic structures into the LLMs to assess their impact on the models’ effectiveness in detecting vulnerabilities.
@@ I-A Research Questions
RQ1. How do various pre-trained LLMs and traditional deep learning techniques perform in detecting security vulnerabilities of Solidity smart contracts?
RQ2. How does fine-tuning of pre-trained LLMs using our proposed domain-specific dataset, VulSmart, impact the performance of vulnerability detection in smart contracts?
RQ3. How do different prompting techniques influence the performance of both pre-trained and fine-tuned LLMs in vulnerability detection?
RQ4: To what extent do adversarial attacks affect the vulnerability detection capabilities of fine-tuned LLMs?
RQ5: Can the integration of semantic structures, such as control-flow and data dependencies, enhance the performance of LLMs in identifying vulnerabilities?
@@ I-B Contributions
  • We plan to make datasets, source codes and models accessible for further research at https://anonymous.4open.science/r/DetectionMadeEasy/.
@@ II Related Studies
The following work have been relevant to the following research areas, namely smart contract vulnerability detection using: a) Formal method tools, and b) Large Language Models (LLMs)
@@ II-A Formal Method Tools
II-B Large Language Models
@@ V-D Observations
RQ1. How do various pre-trained LLMs and traditional deep learning techniques perform in detecting security vulnerabilities of Solidity smart contracts?
@@
RQ2. How does fine-tuning of pre-trained LLMs using our proposed domain-specific dataset, VulSmart, impact the performance of vulnerability detection in smart contracts?
@@
RQ3. How do different prompting techniques influence the performance of both pre-trained and fine-tuned LLMs in vulnerability detection?
@@
[…]
Figure 5: MCC Scores for Vulnerability Classification using Few-Shot across LLMs
Llama2CodeLlamaCodeT5FalconGPT-3.5 TurboGPT-4o Mini000.20.20.20.20.40.40.40.40.60.60.60.60.80.80.80.811110.110.110.110.110.150.150.150.150.290.290.290.290.230.230.230.230.660.660.660.660.710.710.710.71000.170.170.170.170.760.760.760.760.830.830.830.83ModelsValueBase ModelFine-Tuned Model
Figure 6: MCC Scores for Vulnerability Types Class using Few-shot across LLMs
[…]
Figure 7: MCC Scores for Severity Class using Few-shot across LLMs
Llama2CodeLlamaCodeT5FalconGPT-3.5 TurboGPT-4o Mini000.20.20.20.20.40.40.40.40.60.60.60.60.80.80.80.81111000.120.120.120.120.250.250.250.25000.480.480.480.480.530.530.530.530.460.460.460.460.720.720.720.720.250.250.250.250.250.250.250.250.770.770.770.770.960.960.960.96ModelsValueBase ModelFine-Tuned Model
Figure 8: MCC Scores for Vulnerability Classification using Chain-of-thoughts across LLMs
[…]
Figure 9: MCC Scores for Vulnerability Types Class using Chain-of-thoughts across LLMs
[…]
Figure 10: MCC Scores for Severity Class using Chain-of-thoughts across LLMs
Table IV presents the Matthews Correlation Coefficient (MCC) scores, providing insight into the models’ predictive accuracy before and after fine-tuning for both binary classification of vulnerability class and multi-class classification of vulnerability type and severity under zero-shot prompting. Fine-tuning notably improved the performance of all LLMs—open-source and closed-source—except for CodeT5, which maintained an MCC score of 0, indicating it generates random predictions rather than meaningful ones. In contrast, GPT models achieved an MCC score close to 0.99, demonstrating consistently accurate predictions.
The MCC scores for vulnerability detection, vulnerability type, and severity across all models using few-shot prompting are illustrated in Figures 5, 6, and 7, respectively. Likewise, Figures 8, 9, and 10 display the MCC scores for the same tasks under chain-of-thought prompting. All models were fine-tuned using zero-shot prompting, yet we observe that fine-tuned CodeLlama outperformed other open-source LLMs across all prompting techniques, showing strong adaptability to the vulnerability detection tasks.
RQ4: To what extent do adversarial attacks affect the vulnerability detection capabilities of fine-tuned LLMs?
@@
RQ5: Can the integration of semantic structures, such as control-flow and data dependencies, enhance the performance of LLMs in identifying vulnerabilities?
TABLE XII: Performance Comparison of GPT-4o Mini Across Various Semantic Structures
Metrics (in %)
[…]
Only Code Vulnerability 71 61 71 66
Vulnerability Type 24 39 24 23
Severity 41 22 41 28
Code + Control Flow Graph (CFG) Vulnerability 71 62 71 66
Vulnerability Type 24 39 24 23
Severity 44 37 44 40
Code + Data Dependency information Vulnerability 74 62 74 67
Vulnerability Type 26 43 26 24
Severity 44 47 44 37
Code + CFG + Data Dependency Vulnerability 79 81 79 80
Vulnerability Type 29 52 29 28
Severity 47 62 47 45
To assess the impact of integrating semantic structures with Solidity code on the performance of large language models (LLMs), we augment the input prompt to the model with Control Flow Graph (CFG) and data dependency information. Specifically, we evaluate the performance of the GPT-4o Mini model by incorporating these additional semantic features. We select GPT-4o Mini for this evaluation due to its token limit capabilities, which allow for processing larger inputs without information loss. This is critical, as Solidity code combined with semantic structures like CFG and data dependencies often involves a significant number of tokens. With a maximum context window of 128k tokens, GPT-4o Mini is well-suited to handle these extensive inputs, ensuring comprehensive analysis during vulnerability detection in Solidity code.
For our experiment, we select 50 smart contracts, both vulnerable and non-vulnerable, whose CFG and data dependency information could be extracted using the Slither tool [20]. Slither generates the control flow graph (CFG) for each function within a smart contract; however, it often encounters difficulties in generating complete CFGs or providing accurate data dependency information for many real-world smart contracts. The CFG information generated by Slither is in digraph dot language format, which we convert into a paired mapping representation, CFG(Stmt1, Stmt2). This representation indicates that statement Stmt2 is control-flow dependent on statement Stmt1 for the given code statements (i.e. Stmt2 is child of Stmt1 within the control flow of the given code). The data dependency information for each function is provided in the form of key-value pairs, DD(v,V) where the v represents a unique variable, and the V is a list of variables on which v is dependent. To evaluate the effectiveness of GPT-4o Mini in detecting vulnerabilities, their types, and severity levels, we design four different categories of zero-shot prompts: (a) Solidity code alone, (b) Solidity code with its corresponding CFG, (c) Solidity code with its data dependency information, and (d) Solidity code with both CFG and data dependency information.
The performance of GPT-4o Mini under these various settings is summarized in Table XII, using metrics such as accuracy, precision, recall, and F1-score. Our analysis reveals that for binary classification of vulnerability presence, as well as multi-class classification of vulnerability type, incorporating CFG (Control Flow Graph) alongside the code in the prompts does not lead to noticeable improvements over using code alone. However, when it comes to multi-class classification of vulnerability severity, the inclusion of CFG with the code results in an accuracy increase from 41% to 44%, along with improvements in precision, recall, and F1-score compared to prompts using only the code. However, when data dependency information is added to the code, we observe a significant improvement in performance—accuracy increased from 71 to 74, precision increased from 61 to 62, and the F1-score improved from 66 to 67 for vulnerability class. Similarly, incorporating both CFG and data dependency information along with the code led to improved performance compared to using code, code with CFG and code with data dependency information.
These results suggest that control flow information for each function does not contribute significant additional value to vulnerability detection. In contrast, leveraging data dependency information, both independently and in combination with control flow graph (CFG) information, significantly enhances the model’s understanding of the underlying semantics, leading to better detection of vulnerabilities. The data dependency provides richer semantic insights for vulnerability detection compared to control-flow information. Additionally, we conclude that incorporating control flow information alongside data dependency significantly enhances model performance compared to using data dependency information alone.
@@ VII Conclusion and future scope
This paper presents a comprehensive approach to leveraging the power of large language models (LLMs) for detecting vulnerabilities in Solidity-based smart contracts. While much of the existing research highlights the potential of closed-source GPT models and Llama2, we introduce the Smart Contract Vulnerability Detection Framework (SmartVD), which demonstrates superior performance in vulnerability detection. Our findings indicate that among open-source LLMs, the fine-tuned CodeLlama (SmartVD) outperforms and along with it perform better than closed-source base models GPT-3.5 Turbo and GPT-4o Mini. On the other hand, after fine-tuning, the closed-source GPT-3.5 Turbo and GPT-4o Mini models achieve the highest performance across both open and closed-source LLMs.
@@
Although this work adapts multi-class classification based on the type of vulnerabilities present in the code and their severity levels, our future plan is to strengthen this by pinpointing the exact line or vulnerable function within the code. As highlighted in RQ5, incorporating data dependency in combination with control-flow information significantly improves the model’s understanding of the underlying semantics, resulting in enhanced vulnerability detection. In future work, we plan to further explore the impact of integrating richer semantic information into LLM learning. This will include the use of Augmented Control Flow Graphs (CFG), Program Dependency Graphs (PDG), System Dependency Graphs (SDG), and pattern based code slicing information, alongside the source code. This integration has significant potential for advancing vulnerability detection tasks.